OpenStudy (anonymous):

sorry, can someone please solve this for me! i keep getting a weird answer after i used the quadratic formula. solution set for the equation -2x^2+8x-8=0 A) {-2} B) {2} C) {-2,2} D) {-2, 0, 5}

OpenStudy (anonymous):

what's your "wierd" answer?

OpenStudy (anonymous):

or maybe im solvign ti wrong because i get a negative answer inside the radical =P

OpenStudy (unklerhaukus):

\[x=\frac{-8\pm \sqrt{(-8)^2-4(-2)(-8)}}{2(-2)}\]

OpenStudy (anonymous):

that's prolly it... you can also factor out a -2 first to simplify the equation.

OpenStudy (unklerhaukus):

\[\sqrt{(-8)^2-4(-2)(-8)}=64-64=0\]

OpenStudy (unklerhaukus):

*\[=\sqrt{64-64}=\sqrt0=0\] the discriminant is zero

OpenStudy (unklerhaukus):

so \[x=\frac{-8}{-4}=2\]

OpenStudy (anonymous):

alright, got it. thanks!